Under water, dolphins emit sounds which have a wavelength of 5.76 mm. In air the same sounds have a wavelength of 1.36 mm. From this information, and the fact that the speed of sound is 343 m/s in air, how large is the speed of sound in water?
Answer #1

as we know speed of sound

v = f*lemda

AIR ) f = v/lemda = 343/(1.36*10^-3) = 252.20*10^3 Hz

WATER) v = f * lemda = 252.20*10^3*5.76*10^-3 = 1452.70 m/s